OpenStudy (anonymous):

Find the volume of a square pyramid with base 12 inches by 12 inches, height 7.8 inches and slant height 11 inches.

OpenStudy (anonymous):

144 * 7.8 = 1123.3/3 = ?

OpenStudy (anonymous):

1123.2

OpenStudy (anonymous):

347.43

OpenStudy (anonymous):

@AlexandervonHumboldt2